# Evaluating Long Range Dependency Handling in Code Generation Models using Multi-Step Key Retrieval

### v1

@@ 3.5.3 Effect of Forward References
Figure 6: Effect of forward references by task variant, StarCoderBase-7B with 5 distractors. See Figure 9 in the appendix for all models and number of distractors.
We define a forward reference as a function calling a not-yet-defined function. We find that forward references have a strong negative impact on task performance. Figure 6 shows the performance of StarCoderBase-7B with respect to the number of forward references. Accuracy@3 score drops from 87% to 34% for two-step task at 4k context size when the order of the two task snippets is reversed. We find similar behavior with respect to forward references across all the models in our experiment (Figure 9 in Appendix A).
@@ 3.5.4 Effect of Task Snippet Spread
Figure 7: Effect of spread (normalized to the number of tokens in the context) for StarCoderBase-7B with 4k context size and 5 distractors. For data on other models and context sizes, see Figure 10, Figure 11, and Figure 12.
@@
Figure 7 shows the effect of task snippet spread with StarCoderBase-7B with 4k context length. For the two-step task, as spread increases, the performance gap between zero and one forward reference grows, we find that this result is consistent across models Figure 10. For the three-step task a similar effect is clearly visible at 2k and 4k context size. However this effect is not consistent across different models Figure 11.
In the concatenation task, the effect is less pronounced as model’s performance is low regardless of spread and the number of forward references and there is not much room for separation Figure 12.
@@ 4 Improving Retrieval Performance with Call Graph Comments
We saw in Section 3.5.3 that performance degrades for the multi-step tasks in the presence of forward references. We considered if injecting information about function call relationships into the prompt would help performance. This information is easy to obtain from static analysis of code (or in our case at prompt construction time), and if beneficial, provides a lightweight approach to improving performance at inference time. We focus on our most difficult tasks, namely the three-step and concatenation retrieval task with 5 distractors using the same models as in our previous experiment.
@@ 4.1 Experiment Setup
We add comments above each function that contain a lightweight description of the call graph associated with that function. We can construct these comments by annotating which functions are “called by” other functions, or annotating which functions “call” a given function, or a combination of both. In the three-step retrieval task we can do this transitively and add annotation of all the functions that will be called to compute the result of a given function.
We considered two templates for these comments, in one variation (“names only”) we simply included a comma-separated list of function names. In the other (“full sentence”) we use the phrases: “This function is called by ” and “This function calls ” followed by the comma-separated list of function names. See Appendix G for examples.
@@ 4.2 Results
(a) By template
Template Comment Types
Calls Called By Both
Full-Sentence 20.4 31.7 36.7
Names-Only 15.7 28.3 27.2
(b) By task (full sentence template only)
Task Comment Types
None Calls Called By Both
Three-step 32.1 39.1 64.6 68.5
Concatenation 4.5 9.2 11.9 17.7
We observe that the addition of call-graph comments has a positive effect on task performance. 2(b) shows a ~4x improvement on the concatenation task, and a ~2x improvement on the three-step retrieval task with the addition of call-graph comments aggregated across models. We do note that StarCoderBase-1B sees no improvement on the concatenation task. See Appendix B for detailed results.
Full sentence vs. function names only:
The full-sentence template performs better than the names-only version (2(a)). This suggests that the models are able to take some advantage of the more complete natural language description and that there may be room for further improvement by tuning the template. We focus the rest of our analysis on the full-sentence condition.
Call graph directions:
2(b) shows that while most of the benefit comes from the “X is called by Y” type of comment, using both directions produces the best improvement overall.
@@
0 38.0 61.3 67.6
1 33.5 62.9 71.0
2 20.8 47.0 59.3
@@
Figure 8: Effect of call graph comments on StarcoderBase-7B performance across context sizes, spread of task snippets and number of forward references.
@@
Figure 8 shows that call-graph comments improve performance across all context lengths, number of forward references, and task-snippet spreads for StarCoderBase-7B (this trend holds for other models where the call graph comments make a difference). Full plots for all the models are found in Appendix B. We note that with respect to spread for the concatenation task, most of the improvement appears when the task-relevant snippets are closer together.
@@ 5 Discussion
We found that models were able to perform well at the simpler one-step recall task over their entire context lengths (2(a)). However our experiments highlight the importance of distinguishing between long-context support and the ability to handle long-range dependencies. This is illustrated by models like Mistral and StarCoder2 that appear to make a trade-off between the two. While sliding window attention greatly expands their supported context sizes, our experiments show that when distance between relevant snippets is greater than the sliding window size used by these models, performance degrades quickly. Our tasks require precise recall over a long distance and it appears to be difficult to pass on information from earlier sliding windows in a way that allows successful completion of the task.
As different approaches to scaling attention to larger context windows proliferate, practitioners should take care to understand and mitigate these effects if their task relies on precise recall of long-range dependencies. In the context of code generation, builders of retrieval augmented generation systems may find it desirable to rewrite code to bring code snippets that depend on each other closer together.
Our multi-step tasks test the ability of models to reason through indirection introduced by simple yet common operations like function calling and concatenation. We find that models perform significantly worse on these tasks than the simpler one-step recall task (2(a)). This is particularly true in the presence of distractors, which we believe are important for appropriately pressure testing the ability of models to do the required reasoning (Figure 3).
@@
We note that models we tested struggled even more in the presence of forward references (Figure 6). In this situation all the information in the prompt is the same, but is presented in a different order. Humans and compilers are able to resolve this problem and it would be ideal if models could do the same. While we do not know what causes this behaviour, one hypothesis is that the causal attention used during training favors attending backwards in the context window and propagating information from earlier to later tokens.
We found that adding simple natural-language text annotations that describe the call graph greatly improves performance on the more difficult multi-step tasks in all cases, including where forward references are present. Most of the benefit seems to come from just having the function name of the forward-reference appear first, though there also appears to be room to explore tuning the comment text to improve performance (2(b)).
These findings suggest a couple of avenues for practitioners to explore in constructing prompts for code generation. Firstly it may be possible to rewrite code to reduce forward references while maintaining the semantics of the code, and secondly that extra information that can be gleaned from existing methods such as static analysis can be injected as comments to help models reason over the code.
@@ 6 Limitations
Our experiments uses simple synthetically generated functions that are not as realistic as real code in the wild. Recommendations such as re-ordering functions to bring dependant code closer together or adding annotations like call-graph comments would need to be tested in real retrieval-augmented code generation environments. We do not do in-depth investigation of potential internal mechanisms that lead to the behaviours we observed, we describe some hypotheses but otherwise leave that to future work.
@@ 7 Acknowledgements
We thank Alan Leung, Titus Barik, Dominik Moritz, Samira Abnar, Forrest Huang, Jeff Nichols and Barry-John Theobald for their helpful feedback and support
Appendix A Detailed Results: Experiment 1
@@
Table 4: Overall accuracy@3 scores for each model, task, and context size. Each score in the context length columns is the mean across all three distractor conditions. The “mean” column shows the average score for the three context sizes.
@@
one-step Mistral-7B-v0.1 99.8±0.1 99.8±0.1 46.4±0.9 61.5
StarCoder2-7B 100.0±0.0 99.8±0.1 68.1±0.8 67.0
StarCoderBase-1B 97.1±0.3 98.3±0.2 98.5±0.1 73.5
StarCoderBase-7B 100.0±0.0 99.9±0.0 99.6±0.1 74.9
StarCoderBase-15.5B 99.8±0.1 99.5±0.1 99.0±0.1 74.6
two-step Mistral-7B-v0.1 63.1±0.7 58.3±0.6 16.7±0.5 34.6
StarCoder2-7B 81.9±0.5 82.6±0.5 46.4±0.6 52.7
StarCoderBase-1B 73.3±0.6 73.5±0.5 73.7±0.5 55.1
StarCoderBase-7B 81.6±0.5 80.1±0.5 79.2±0.5 60.2
StarCoderBase-15.5B 82.2±0.5 74.7±0.5 64.0±0.6 55.2
three-step Mistral-7B-v0.1 53.1±0.6 41.5±0.5 14.7±0.4 27.3
StarCoder2-7B 78.9±0.5 75.3±0.5 40.0±0.6 48.6
StarCoderBase-1B 55.9±0.6 55.2±0.6 52.9±0.6 41.0
StarCoderBase-7B 67.0±0.6 66.9±0.5 66.2±0.6 50.0
StarCoderBase-15.5B 65.8±0.6 57.1±0.6 45.3±0.6 42.1
concatenation Mistral-7B-v0.1 17.3±0.4 6.7±0.3 0.7±0.1 6.2
StarCoder2-7B 34.7±0.6 18.3±0.5 4.5±0.3 14.4
StarCoderBase-1B 1.7±0.1 0.6±0.1 0.1±0.0 0.6
StarCoderBase-7B 21.9±0.5 13.4±0.4 8.0±0.3 10.8
StarCoderBase-15.5B 34.4±0.6 21.6±0.5 15.2±0.4 17.8
@@
Context Size 2k 4k 8k # Distractors 0 1 5 0 1 5 0 1 5 Mistral-7B-v0.1 100.0 99.9 99.6 100.0 100.0 99.5 48.5 48.2 42.8 StarCoder2-7B 100.0 100.0 99.9 100.0 100.0 99.6 76.3 69.3 59.8 StarCoderBase-1B 100.0 99.1 93.2 99.9 98.4 96.7 99.9 98.8 97.0 StarCoderBase-7B 100.0 100.0 99.9 100.0 100.0 99.9 100.0 99.7 99.1 StarCoderBase-15.5B 99.9 99.8 99.8 100.0 99.5 99.0 99.8 99.3 98.1
@@
Context Size 2k 4k 8k # Distractors 0 1 5 0 1 5 0 1 5 Mistral-7B-v0.1 87.4 74.5 40.6 76.7 63.9 34.4 22.7 18.9 8.6 StarCoder2-7B 99.2 85.5 65.5 97.9 86.8 63.0 62.6 49.1 27.5 StarCoderBase-1B 98.5 80.9 47.3 96.2 82.5 41.8 89.5 79.2 52.4 StarCoderBase-7B 99.2 86.3 64.0 97.9 81.9 60.6 95.4 85.3 56.8 StarCoderBase-15.5B 93.6 80.9 74.4 89.6 68.8 65.8 82.0 64.3 45.8
@@
Context Size 2k 4k 8k # Distractors 0 1 5 0 1 5 0 1 5 Mistral-7B-v0.1 81.4 54.0 23.8 61.4 42.3 21.0 19.5 14.7 10.0 StarCoder2-7B 95.6 83.0 58.2 94.5 85.4 46.0 60.0 39.5 20.5 StarCoderBase-1B 80.7 59.0 28.1 70.9 68.3 26.4 79.6 58.7 20.5 StarCoderBase-7B 91.7 68.0 41.2 82.6 75.9 42.3 88.0 77.1 33.4 StarCoderBase-15.5B 81.4 66.5 49.5 69.7 62.1 39.6 65.0 49.5 21.5

### v2

@@ 3.5.4 Effect of Forward References
Table 2: Model accuracy (%) vs. number of forward references. 5-distractor condition. 95% confidence interval shown adjacent. Performance generally drops as number of forward references increases. Blue indicates exceptions to this behavior, indicating a condition that performs worse than the condition with one more forward reference.
Two-step Task
Context Size 2k 4k 8k # Forward References 0 1 0 1 0 1 Mistral-7B-v0.1 60.1$\pm$1.1 21.1$\pm$1.2 46.4$\pm$1.3 22.5$\pm$1.3 8.0$\pm$1.3 9.2$\pm$1.3 StarCoder2-7B 73.5$\pm$0.1 57.6$\pm$0.4 69.4$\pm$0.4 56.5$\pm$0.8 28.1$\pm$0.6 26.8$\pm$1.2 StarCoderBase-1B 55.3$\pm$0.4 39.4$\pm$0.7 49.7$\pm$1.0 33.9$\pm$1.1 65.6$\pm$1.1 39.3$\pm$1.3 StarCoderBase-7B 90.0$\pm$1.4 38.1$\pm$1.1 87.2$\pm$1.4 34.1$\pm$1.1 76.0$\pm$0.8 37.7$\pm$0.8 StarCoderBase-15.5B 85.5$\pm$1.3 63.4$\pm$1.4 82.3$\pm$1.4 49.2$\pm$1.4 63.4$\pm$1.4 28.3$\pm$1.3 DeepSeekCoder-6.7B-base 83.7$\pm$1.0 75.2$\pm$1.3 74.4$\pm$1.0 67.7$\pm$1.4 70.2$\pm$1.3 58.5$\pm$1.3 GPT-4o-mini 98.4$\pm$1.4 95.7$\pm$1.4 91.4$\pm$1.4 87.2$\pm$1.3 87.4$\pm$1.4 80.3$\pm$1.4 GPT-4o 99.8$\pm$0.8 98.3$\pm$1.3 98.6$\pm$0.9 93.6$\pm$1.3 96.2$\pm$1.2 83.9$\pm$1.3
Three-step Task
Context Size 2k 4k 8k # Forward References 0 1 2 0 1 2 0 1 2 Mistral-7B-v0.1 28.1$\pm$2.1 24.4$\pm$1.0 16.8$\pm$2.2 19.8$\pm$2.3 22.3$\pm$1.1 16.9$\pm$2.1 2.7$\pm$2.3 11.7$\pm$1.2 10.3$\pm$2.1 StarCoder2-7B 61.5$\pm$1.1 60.5$\pm$0.5 45.7$\pm$1.8 46.2$\pm$1.9 48.1$\pm$0.8 37.5$\pm$2.0 14.4$\pm$2.0 23.1$\pm$0.8 16.0$\pm$1.9 StarCoderBase-1B 32.3$\pm$2.8 29.5$\pm$1.1 18.1$\pm$2.7 29.5$\pm$3.0 28.5$\pm$1.4 14.9$\pm$2.8 19.8$\pm$2.8 23.8$\pm$1.4 8.2$\pm$2.8 StarCoderBase-7B 63.4$\pm$2.0 41.3$\pm$0.9 18.8$\pm$1.7 65.4$\pm$1.8 42.0$\pm$0.9 20.5$\pm$1.6 41.9$\pm$0.8 35.6$\pm$0.8 15.9$\pm$1.3 StarCoderBase-15.5B 57.2$\pm$2.4 50.6$\pm$1.2 37.2$\pm$2.4 54.0$\pm$2.4 39.7$\pm$1.3 24.9$\pm$2.4 34.5$\pm$1.8 21.0$\pm$1.1 10.5$\pm$1.7 DeepSeekCoder-6.7B-base 69.4$\pm$2.4 73.1$\pm$1.2 56.2$\pm$2.3 52.2$\pm$2.3 61.3$\pm$1.2 52.6$\pm$1.9 42.9$\pm$2.2 50.5$\pm$1.0 40.1$\pm$1.3 GPT-4o-mini 68.0$\pm$2.2 81.8$\pm$1.1 70.9$\pm$1.9 53.3$\pm$2.3 67.1$\pm$1.1 62.8$\pm$1.8 39.4$\pm$2.0 56.3$\pm$1.1 50.4$\pm$1.3 GPT-4o 95.7$\pm$2.2 96.7$\pm$1.2 89.6$\pm$1.6 88.7$\pm$2.0 91.3$\pm$1.1 86.6$\pm$1.8 85.2$\pm$2.3 91.0$\pm$1.1 86.8$\pm$1.6
Concatenation Task
Context Size 2k 4k 8k # Forward References 0 1 2 0 1 2 0 1 2 Mistral-7B-v0.1 11.8$\pm$1.8 4.3$\pm$1.5 3.1$\pm$1.3 2.4$\pm$1.2 0.6$\pm$0.9 0.6$\pm$0.8 0.8$\pm$1.0 0.2$\pm$0.6 0.2$\pm$0.5 StarCoder2-7B 17.7$\pm$0.4 14.7$\pm$0.6 7.9$\pm$0.6 6.4$\pm$0.6 5.4$\pm$0.7 4.2$\pm$0.8 1.8$\pm$0.8 0.7$\pm$0.9 0.7$\pm$1.2 StarCoderBase-1B 1.6$\pm$1.6 0.3$\pm$1.8 0.0$\pm$1.9 0.5$\pm$1.9 0.1$\pm$2.0 0.0$\pm$2.0 0.0$\pm$2.0 0.0$\pm$1.9 0.0$\pm$1.8 StarCoderBase-7B 13.9$\pm$1.1 9.3$\pm$0.7 4.0$\pm$0.5 2.0$\pm$0.5 1.6$\pm$0.2 1.2$\pm$0.2 0.6$\pm$0.3 0.6$\pm$0.1 0.3$\pm$0.2 StarCoderBase-15.5B 22.9$\pm$1.3 14.2$\pm$1.3 11.0$\pm$0.9 10.1$\pm$0.9 6.1$\pm$0.8 5.8$\pm$0.7 6.8$\pm$0.5 4.5$\pm$0.3 2.6$\pm$0.3 DeepSeekCoder-6.7B-base 40.3$\pm$1.5 24.7$\pm$1.2 20.2$\pm$1.1 14.2$\pm$1.1 8.5$\pm$0.8 7.2$\pm$0.8 9.6$\pm$0.8 4.6$\pm$0.7 3.2$\pm$0.5 GPT-4o-mini 79.8$\pm$0.4 71.2$\pm$0.2 63.6$\pm$0.0 64.9$\pm$0.2 52.7$\pm$0.1 45.5$\pm$0.0 46.1$\pm$0.0 35.4$\pm$0.0 27.7$\pm$0.0 GPT-4o 98.5$\pm$1.2 96.7$\pm$1.0 97.3$\pm$0.6 97.4$\pm$0.5 96.0$\pm$0.4 94.9$\pm$0.3 93.6$\pm$0.2 91.7$\pm$0.2 88.9$\pm$0.1
We define a forward reference as a function calling a not-yet-defined function (i.e. one that will be defined later in the prompt). We find that forward references have a negative impact on task performance in all models, in most conditions we tested. The exceptions to this are highlighted in blue in Table 2, and mainly occur in the three-step task where we sometimes observe an increase in performance when going from zero to one forward reference followed by a drop going from two to three forward references. Table 2 shows these results in detail. We see drops as large as 44.9% in models like StarCoderBase-7B (three-step task at 4k), 19.4% for GPT-4o-mini (concatenation task at 4k) and 12.3% for GPT4o (two-step task at 8k).
@@ 3.5.5 Effect of Task Snippet Spread
StarCoderBase-7B
DeepSeekCoder-6.7B-base
Figure 7: Effect of spread (normalized to the number of tokens in the context) for StarCoderBase-7B (top) and DeepSeekCoder-6.7B-base (bottom) with 4k context size, 5-distractor condition. Shaded regions represent 95% confidence interval. For data on other models and context sizes, see Figure 10, Figure 11, and Figure 12.
@@
In the two-step task we broadly see three behaviors with respect to spread, the GPT4o models show little variation as spread increases. DeepSeekCoder-6.7B-base, StarCoder2-7B and Mistral-7B-v0.1 show performance drops at moderate amount of spreads but perform well when snippets are very close to each other or very far from each other. Snippets with maximal spread are located close to the beginning and end of the prompt and result in higher performance. The other models (StarCoderBase-1B, StarCoderBase-7B, and StarCoderBase-15B) show a drop in performance as spread increases, but only in the presence of forward references. Figure 7 shows the effect of task snippet spread for StarCoderBase-7B DeepSeekCoder-6.7B-base with 4k context length on the two-step retrieval task with results for the other models in the appendix (Figure 10).
For the three-step task similar, effects are clearly visible at 2k and 4k context size but are not as consistent across different models Figure 11. For the concatenation task, the effect is less pronounced as model’s performance is generally low regardless of spread and the number of forward references and there is not much room for separation Figure 12.
@@ 4 Improving Retrieval Performance with Call Graph Comments
We saw in Section 3.5.4 that performance degrades for the multi-step tasks in the presence of forward references. Loosely inspired by chain-of-thought prompting (Wei et al., 2023; Kojima et al., 2023) that encourages models to produce intermediate token representing sub-parts of the problem, we considered if injecting information about function call relationships into the prompt would help performance. This allows us to control the amount of additional information about the problem provided and importantly, where it appears. So unlike chain-of-thought prompting, we do not allow the model to produce extra tokens but rather inject ‘extra’ tokens into particular areas of the prompt.
We focused on our most difficult tasks, namely the three-step and concatenation retrieval task with 5 distractors using the same models as in our previous experiment.
@@ 4.1 Experiment Setup
We add comments above each function that contain a lightweight description of the call graph associated with that function. We can construct these comments by annotating which functions are “called by” other functions, or annotating which functions “call” a given function, or a combination of both. In the three-step retrieval task we can also do this transitively and add annotation of all the functions that will be called to compute the result of a given function (as opposed to just the next function in the chain).
We considered two templates for these comments, in one variation (“names only”) we simply included a comma-separated list of function names and no other information. This has the effect of introducing the tokens for the referenced function before it appears as a call target, though without any indication that it is even a function name.
In the other template (“full sentence”) we use the phrases: “This function is called by ” and “This function calls ” followed by the comma-separated list of function names. See Appendix I for examples.
@@ 4.2 Results
Comment Type
Template none calls called-by both
Three-step Names Only 46.9 47.8 72.9 69.3
Full Sentence 46.9 55.3 74.2 76.8
Concatenation Names Only 23.3 24.6 29.5 29.5
Full Sentence 23.3 27.6 32.8 36.4
We observe that the addition of call-graph comments has a positive effect on task performance. Table 3 shows a ~1.5x improvement on the concatenation task, and a ~1.6x improvement on the three-step retrieval task with the addition of call-graph comments aggregated across models. We do note that StarCoderBase-1B sees no improvement on the concatenation task. See Appendix D for detailed results.
@@
0 48.6 70.4 73.9
1 48.8 71.9 79.0
2 37.8 58.9 71.0
Full sentence vs. function names only:
The full-sentence template performs better than the names-only version (Table 3). This suggests that the models are able to take some advantage of the more complete natural language description and that there may be room for further improvement by tuning the template. However it is notable that simply including the names of the referenced functions before their function definition has such a large performance boost. We focus the rest of our analysis on the full-sentence condition as it was the best performing template.
Call graph directions:
Table 3 shows that while most of the benefit comes from the “X is called by Y” type of comment, using both directions produces the best improvement overall.
@@
StarCoderBase-7B
GPT-4o-mini
Figure 8: Effect of call graph comments on StarcoderBase-7B (top) and GPT-4o-mini (bottom) performance across context sizes, spread of task snippets and number of forward references. Shaded regions represent 95% confidence interva
@@
Figure 8 shows that call-graph comments improve performance across all context lengths, number of forward references, and task-snippet spreads for StarCoderBase-7B (this trend holds for other models where the call graph comments make a difference). Full plots for all the models are found in Appendix D. We note that with respect to spread for the concatenation task, most of the improvement appears when the task-relevant snippets are closer together.
@@ 5 Discussion
Our findings align with those of Yu et al. (2024), who show that over 70% of functions in popular open-source repositories are “non-standalone” and report a 48% average performance drop when models are evaluated on such functions. Their report underscores the need to evaluate the ability of models to reason over interdependent code units. Our use of synthetic code enables us to expand the context beyond the 1K token limit in their benchmark and gain fine-grain control over the relations being tested.
Our results also help distinguish between long-context support and the ability to handle long-range dependencies. We find that models were able to perform well at the simpler one-step recall task over their entire context lengths (2(a)), but performance degrades as the number of steps required increases. This is further exacerbated by the presence of distractors which we believe are important for appropriately pressure testing the ability of models to do the required reasoning (Figure 3).
@@
We also observe architectural choices that appear to make a tradeoff between efficient support of long-context inputs and capturing long-range token relationships. This is illustrated by models like Mistral and StarCoder2, whose use of sliding window attention greatly expands their supported context sizes, however our experiments show that when distance between relevant snippets is greater than the sliding window size used by these models, performance degrades quickly. Our tasks require precise recall over a long distance and it appears to be difficult to pass on information from earlier sliding windows in a way that allows successful completion of the task. As different approaches to scaling attention to larger context windows proliferate, practitioners should take care to understand and mitigate these effects if their task relies on precise recall of long-range dependencies. In the context of code generation, builders of retrieval augmented generation systems may find it desirable to rewrite code to bring code snippets that depend on each other closer together.
Our findings shed light on specific factors influencing the fragility of multi-step reasoning in large language models, including the presence of distractors, function ordering and architectural constraints such as sliding-window attention.
@@
We note that models we tested struggled even more in the presence of forward references (Table 2). In this situation all the information in the prompt is the same, but is presented in a different order. Humans and compilers are able to resolve this problem and it would be ideal if models could do the same. While we do not know what causes this behavior, one hypothesis is that the causal attention used during training favors attending backwards in the context window and propagating information from earlier to later tokens.
While we leave the discovery of the mechanisms that underlie such behavior to future work, we do observe useful progress in the field that could be brought to bear. While there is still debate on whether examining attention patterns in model can provide faithful explanations (Jain & Wallace, 2019; Wiegreffe & Pinter, 2019), recent progress in mechanistic interpretability is developing tools to closely examine components involved in resolving relationships between tokens and the mechanisms transformers use to encode them (e.g. Elhage et al. (2021)). Hernandez et al. (2024) find that a subset of natural-language entity relations can be modeled with a single linear transformation. Brinkmann et al. (2024) analyze mechanisms involved in symbolic multi-step reasoning in transformers using a synthetic tree-traversal task and describe mechanism such as ”backward-chaining“ and ”one-step“ lookahead in the models they train. Such techniques could be applied to larger pre-trained models to understand the failure modes in resolving forward references that we observed in this work.
Finally we find that we can add information to the prompt that may short-circuit whichever mechanisms are involved in resolving function references. We see that introducing just the tokens of referenced functions before they are used in a function definition greatly improves performance on the more difficult multi-step tasks in all cases, especially where forward references are present (Table 3). We believe these scenarios provide a rich testbed to better examine how information flows across tokens particularly in scenarios where models must ‘plan’ for the future.
@@ 6 Limitations
Our experiments uses synthetically generated functions that are not as realistic as real code in the wild. Our tests only use top level functions composed via function call chains and concatenation. Nor do we test more complex structures of indirection such as recursion, branching control flow, or abstractions that cause functions in the call graph to be reused (e.g. helper functions used by multiple functions). These may introduce additional retrieval challenges or sensitivity to ordering not captured by our current setup. Furthermore our use of randomly generated function names results in a setting that is likely out-of-domain. Models may perform better with more realistic identifiers drawn from natural codebases.
We also evaluate only decoder-only causal transformers. Models with recurrence or bidirectional attention may behave differently and merit separate study, particularly with regard to the forward reference finding.
Suggestions such as re-ordering functions to bring dependent code closer together or adding annotations like call-graph comments would need to be tested in real retrieval-augmented code generation environments and should be seen as exploratory; we are not recommending these approaches over others that may improve performance on these tasks (e.g. fine tuning).
@@ 7 Acknowledgements
We thank Alan Leung, Titus Barik, Dominik Moritz, Samira Abnar, Forrest Huang, Jeff Nichols and Barry-John Theobald for their helpful feedback and support.
Appendix A Appendix
Appendix B Detailed Results: Experiment 1
@@
Table 5: Overall accuracy@3 scores for each model, task, and context size. Each score in the context length columns is the mean across all number of distractors. The “mean” column shows the average score for the three context sizes.
@@
one-step Mistral-7B-v0.1 99.8±0.6 99.8±0.6 46.4±0.5 82.0
StarCoder2-7B 100.0±0.2 99.8±0.3 68.1±0.3 89.3
StarCoderBase-1B 97.1±0.5 98.3±0.6 98.5±0.7 98.0
StarCoderBase-7B 100.0±0.4 99.9±0.3 99.6±0.1 99.8
StarCoderBase-15.5B 99.8±0.6 99.5±0.5 99.0±0.3 99.5
DeepSeekCoder-6.7B-base 99.7±0.6 99.2±0.5 98.8±0.4 99.2
GPT-4o-mini 99.9±0.1 99.8±0.1 99.3±0.0 99.6
GPT-4o 100.0±0.5 99.9±0.4 99.6±0.3 99.9
two-step Mistral-7B-v0.1 63.1±0.1 58.3±0.2 16.7±0.2 46.1
StarCoder2-7B 81.9±0.0 82.6±0.1 46.4±0.1 70.3
StarCoderBase-1B 73.3±0.1 73.5±0.1 73.7±0.1 73.5
StarCoderBase-7B 81.6±0.1 80.1±0.1 79.2±0.9 80.3
StarCoderBase-15.5B 82.2±0.0 74.7±0.1 64.0±0.8 73.7
DeepSeekCoder-6.7B-base 90.1±0.1 88.0±0.1 85.6±0.1 87.9
GPT-4o-mini 98.8±0.3 95.8±0.2 92.2±0.1 95.6
GPT-4o 99.1±0.0 97.9±0.0 95.5±0.1 97.5
three-step Mistral-7B-v0.1 53.1±0.4 41.5±0.4 14.7±0.5 36.5
StarCoder2-7B 78.9±0.2 75.3±0.3 40.0±0.3 64.7
StarCoderBase-1B 55.9±0.4 55.2±0.5 52.9±0.6 54.7
StarCoderBase-7B 67.0±0.6 66.9±0.5 66.2±0.4 66.7
StarCoderBase-15.5B 65.8±0.5 57.1±0.5 45.3±0.6 56.1
DeepSeekCoder-6.7B-base 88.3±0.6 83.9±0.6 79.0±0.6 83.7
GPT-4o-mini 90.8±0.6 86.5±0.6 77.8±0.6 85.0
GPT-4o 97.9±0.6 95.6±0.5 94.1±0.6 95.9
concatenation Mistral-7B-v0.1 17.3±0.4 6.7±0.4 0.7±0.4 8.2
StarCoder2-7B 34.7±0.1 18.3±0.2 4.5±0.3 19.2
StarCoderBase-1B 1.7±0.2 0.6±0.3 0.1±0.4 0.8
StarCoderBase-7B 21.9±0.7 13.4±0.6 8.0±0.5 14.4
StarCoderBase-15.5B 34.4±0.5 21.6±0.5 15.2±0.6 23.7
DeepSeekCoder-6.7B-base 46.2±0.5 32.5±0.5 24.6±0.6 34.4
GPT-4o-mini 78.6±0.6 70.9±0.5 57.2±0.5 68.9
GPT-4o 97.9±0.5 95.5±0.5 92.9±0.5 95.4
@@
Context Size 2k 4k 8k # Distractors 0 1 5 0 1 5 0 1 5 Mistral-7B-v0.1 100.0 99.9 99.6 100.0 100.0 99.5 48.5 48.2 42.8 StarCoder2-7B 100.0 100.0 99.9 100.0 100.0 99.6 76.3 69.3 59.8 StarCoderBase-1B 100.0 99.1 93.2 99.9 98.4 96.7 99.9 98.8 97.0 StarCoderBase-7B 100.0 100.0 99.9 100.0 100.0 99.9 100.0 99.7 99.1 StarCoderBase-15.5B 99.9 99.8 99.8 100.0 99.5 99.0 99.8 99.3 98.1 DeepSeekCoder-6.7B-base 100.0 100.0 99.3 100.0 100.0 98.0 100.0 99.7 96.8 GPT-4o-mini 100.0 99.6 100.0 100.0 99.7 99.7 100.0 98.7 99.2 GPT-4o 100.0 100.0 100.0 100.0 100.0 99.9 99.9 99.7 99.4
@@
Context Size 2k 4k 8k # Distractors 0 1 5 0 1 5 0 1 5 Mistral-7B-v0.1 87.4 74.5 40.6 76.7 63.9 34.4 22.7 18.9 8.6 StarCoder2-7B 99.2 85.5 65.5 97.9 86.8 63.0 62.6 49.1 27.5 StarCoderBase-1B 98.5 80.9 47.3 96.2 82.5 41.8 89.5 79.2 52.4 StarCoderBase-7B 99.2 86.3 64.0 97.9 81.9 60.6 95.4 85.3 56.8 StarCoderBase-15.5B 93.6 80.9 74.4 89.6 68.8 65.8 82.0 64.3 45.8 DeepSeekCoder-6.7B-base 100.0 96.5 79.4 100.0 92.9 71.0 100.0 92.6 64.4 GPT-4o-mini 100.0 99.4 97.1 100.0 98.1 89.3 100.0 92.7 83.9 GPT-4o 100.0 98.2 99.1 100.0 97.6 96.1 99.0 97.4 90.0
@@
Context Size 2k 4k 8k # Distractors 0 1 5 0 1 5 0 1 5 Mistral-7B-v0.1 81.4 54.0 23.8 61.4 42.3 21.0 19.5 14.7 10.0 StarCoder2-7B 95.6 83.0 58.2 94.5 85.4 46.0 60.0 39.5 20.5 StarCoderBase-1B 80.7 59.0 28.1 70.9 68.3 26.4 79.6 58.7 20.5 StarCoderBase-7B 91.7 68.0 41.2 82.6 75.9 42.3 88.0 77.1 33.4 StarCoderBase-15.5B 81.4 66.5 49.5 69.7 62.1 39.6 65.0 49.5 21.5 DeepSeekCoder-6.7B-base 100.0 95.3 69.7 99.9 93.6 58.3 99.3 90.0 47.5 GPT-4o-mini 100.0 94.8 77.7 99.8 95.7 64.1 98.5 82.3 52.5 GPT-4o 100.0 98.4 95.3 100.0 96.7 90.1 99.8 93.1 89.4